OpenStudy (anonymous):

How many molecules (or formula units) are in each sample? 97.45gCCl4 70.69kgNaHCO3 118.00gC4H10 3.53×104gNa3PO4

OpenStudy (abb0t):

You must use dimensional analysis and the relationship of avogadro's number to find the molecules: Avogadro's number is: \(\large 6.022 \times 10^{23} \frac{molecules}{mole}\) First, find the molar mass of each element. Then find moles of each element. Then once you have moles, use avogadro's number to find the molecules present. If you need more help, let me know. Good luck.

OpenStudy (anonymous):

im still having trouble

OpenStudy (abb0t):

Ok, where are you stuck? I can walk you through this :)

OpenStudy (abb0t):

For the first one, start by calculating the molar mass. What that means is basically to add up the total of each element. In your given problem, you have carbon tetrachloride (also known as benziform) So first, find the mass of each individual elemet (look at your periodic table) carbon = 12.01 chlorine = 35.45 however, notice that you have \(\large 4\) chlorine molecules, which means you have to multiply by 4! so mass of ONE chlorine atom \(35.45 \times 4\) = mass of \(Cl_4\) Now, find the sum. The sum is your molar mass and is in units of \(\large \frac{grams}{mol}\) You can do this by adding \(12.01 + (35.45 \times 4) = molar~ mass \large[\frac{g}{mol}]\) Now, you have to use dimensional analysis. Similar to how you did in algebra, you are cross cnceling out grams using the given mass to simply get mole by itself. Then Use avogadro's number to cancel out mols and you get your molecules. You should get 0.63361508452 moles! or \(6.33 \times 10^{-1}\) moles now cancel out moles and divide by avogadro's number Do the same with the others. Goodluck